two angles are complementary if the sum of their measures is 90° three angles are complementary. one angle is double the smallest angle. the largest angle is triple the smallest angle.
Answer by mananth(16946)


three angles are complementary. one angle is double the smallest angle. the largest angle is triple the smallest angle.
Let smallest angle be x
the second is twice soit is 2x
and the third angle is 3x
x+2x+3x = 90
6x =90
x = 90/6 = 15
(
The angles are 15 ,( 15*2 = 30),( 15*3 = 45) degrees
FIG below
.